OpenStudy (anonymous):

One morning, Justin drove 3 hours before stopping to eat lunch. After lunch, he increased his speed by 10 mph. If he completed a 390-mile trip in 6 hours of driving time, how fast (in miles per hour) did he drive in the morning?

OpenStudy (boholmission):

Speed time formula speed * time = Distance speed = x time1 = 3 hours before lunch time2 = 3 hours after lunch Distance = 390 (time1*3)+(time2*3)=Distance Before lunch speed = x for 3 hours = 3x after lunch speed = x + 10 for 3 hours = 3(x+10) This totals the 6 hours 3x + 3(x+10) = 390 3x + 3x + 30 = 390 6x = 360 x = 60 In the morning he drove 60 MPH check answer In the morning he drove 60 mph for 3 hours totaling 180 miles in the afternoon he drove 70 mph for 3 hours totaling 210 miles 180+210 = 390 miles for the whole trip Correct